# Exploring the Alignment Landscape: LLMs and Geometric Deep Models in Protein Representation

### v1

@@
Abstract
Latent representation alignment has become a foundational technique for constructing multimodal large language models (MLLM) by mapping embeddings from different modalities into a shared space, often aligned with the embedding space of large language models (LLMs) to enable effective cross-modal understanding. While preliminary protein-focused MLLMs have emerged, they have predominantly relied on heuristic approaches, lacking a fundamental understanding of optimal alignment practices across representations. In this study, we explore the alignment of multimodal representations between LLMs and Geometric Deep Models (GDMs) in the protein domain. We comprehensively evaluate three state-of-the-art LLMs (Gemma2-2B, LLaMa3.1-8B, and LLaMa3.1-70B) with four protein-specialized GDMs (GearNet, GVP, ScanNet, GAT). Our work examines alignment factors from both model and protein perspectives, identifying challenges in current alignment methodologies and proposing strategies to improve the alignment process. Our key findings reveal that GDMs incorporating both graph and 3D structural information align better with LLMs, larger LLMs demonstrate improved alignment capabilities, and protein rarity significantly impacts alignment performance. We also find that increasing GDM embedding dimensions, using two-layer projection heads, and fine-tuning LLMs on protein-specific data substantially enhance alignment quality. These strategies offer potential enhancements to the performance of protein-related multimodal models. Our code and data are available at https://github.com/Tizzzzy/LLM-GDM-alignment.
1 Introduction
@@
Recently, multimodal large language models (MLLM) such as GPT-4V [21], Gemini [23], Llama-3.2 [17], LLaVA-1.5 [15], etc., have received increasing attention from the community. A MLLM is an LLM-based model with the ability to receive, reason, and output with multimodal information. This extends beyond just text to potentially include various forms of input and output such as image, audio, video, and other sensory data, allowing the model to process and generate various types of information in an integrated manner. In the realm of protein science, specialized MLLMs such as ProtChatGPT [27], ProteinChat [6] and ProteinGPT [29] have emerged as powerful tools for comprehensive protein analysis. These models can process both protein sequences and structures, integrating this information with vast language understanding capabilities. By combining protein sequence and structure encoders with large language models, these systems can perform complex tasks such as property prediction, structure understanding, and even engage in interactive conversations about protein characteristics. This fusion of modalities allows for a more comprehensive and efficient approach to protein analysis, potentially revolutionizing fields such as drug development and biotechnological advancements.
@@
In these protein-focused MLLMs, latent representation alignment serves as a fundamental technique for integrating different types of protein data. A protein can be represented in multiple modalities - as textual descriptions capturing its function and properties, as graph structures representing amino acid relationships, or as 3D coordinate data describing its spatial configuration. While specialized models exist for each of these representations (like LLMs for text and GDMs for structures), the challenge lies in bridging the gap between these different protein representations to enable comprehensive protein analysis. This is where latent representation alignment plays a crucial role, allowing these diverse protein representations to be mapped into a shared space where they can be meaningfully compared and integrated [27, 6, 29].
@@
Several techniques are commonly used to build multimodal models, each with distinct methods for integrating information from different modalities. In the bio domain, recent models such as Prot2Text [1] integrates Graph Neural Networks (GNNs) [28] and LLMs using an encoder-decoder architecture to generate detailed free-text descriptions of protein functions. MM-StackEns model [2] uses a combination of sequence data processed by a siamese neural network [12] and graph data processed by Graph Attention Networks (GATs) [26] to perform Protein-Protein Interaction (PPI) prediction [11]. ProtST framework [30] uses multimodal representation alignment and multimodal mask prediction tasks to align protein sequence representations with biomedical text representations. Among these methods, latent representation alignment remains the most popular due to its simplicity and scalability [3]. Protein-focused MLLMs such as ProteinGPT [29] is a multimodal chat system designed for protein analysis, combining protein sequence and structure data with a LLM. This model integrates a protein sequence encoder and structure encoder with a linear projection layer to align diverse embeddings with language model inputs. ProteinChat [6] introduces a ChatGPT-like interface for 3D protein structures, emphasizing interactive protein analysis. Its architecture consists of a GVP [10] block to encode structural information, a projection layer, and a LLM. The protein embeddings from the GVP are aligned with the LLM’s language representations through the projection layer, allowing detailed question-based responses regarding protein 3D structures. ProtChatGPT [27] introduces a multimodal LLM framework that combines protein sequence and structure data using a specialized Protein-Language Pertaining Transformer (PLP-former) to bridge the gap between protein structures and text data. The PLP-former aligns protein data with textual descriptions, while a projection adapter facilitates multimodal integration with the LLM, allowing for accurate question-based protein analysis. By aligning the latent spaces of different modalities, this technique enables seamless cross-modal integration without requiring complex fusion mechanisms or retraining weights for every feature.
@@
Despite the current progress in protein-focused MLLMs, latent representation alignment across different protein modalities remains a significant challenge that requires deeper investigation [29]. We aim to address these gaps in understanding and provide insights that could lead to more principled and effective approaches in multimodal protein analysis. In the protein domain, LLMs excel at processing and understanding textual descriptions of proteins, including their functions, properties, and relationships documented in scientific literature [13, 16, 8, 5]. Meanwhile, GDMs are specialized neural architectures designed to capture the complex three-dimensional structural information of proteins, including atomic coordinates, bond angles, and spatial relationships between amino acids [14, 18, 7, 33]. While both model types offer complementary strengths in protein analysis, effectively aligning their representations remains a significant challenge. Our work explores this alignment challenge through three main areas: I) The model perspective analysis, focusing on alignment performance and correlations between different model pairs. II) The protein perspective analysis, examining which types of proteins align well or poorly across models. III) Strategies for improving alignment, including the impact of GDM representation dimensionality, the effect of projection head complexity, and the benefits of fine-tuning LLMs on protein-specific data. To investigate these areas, we conduct extensive experiments comparing three state-of-the-art LLMs (Gemma2-2B, LLaMa3.1-8B, and LLaMa3.1-70B) with four protein-specialized GDMs (GearNet, GVP, ScanNet, GAT). We analyze the alignment performance of different model pairs, study how protein characteristics affect alignment quality, and evaluate various strategies for enhancing alignment through model architecture modifications and training approaches. An overview of our methodology is provided in Figure 1.
@@
Our exploration of the alignment between LLMs and GDMs in the protein domain has revealed several key findings. We found that GDMs integrating both graph and 3D structural information of proteins tend to align better with LLMs. Additionally, larger LLMs with higher embedding dimensions showed improved alignment performance with the same GDM. Our analysis revealed strong correlations between high-alignment model pairs and other high-alignment pairs. Notably, the rarity of a protein significantly affects the model’s alignment performance, with rare proteins posing greater challenges. Based on our observations, we highlight the challenges and limitations of current protein datasets for representation alignment, particularly due to unequal levels of study across proteins and the presence of homologous relationships among them. In terms of model architecture, we discovered that retraining GDMs to have higher embedding dimensions enhances alignment with LLMs. The complexity of the projection head also plays a crucial role, as increasing the number of layers improves alignment up to a certain threshold, beyond which benefits diminish. Finally, we found that fine-tuning LLMs with protein-specific data can significantly improve alignment with GDMs. By revealing the factors that influence alignment performance and suggesting strategies for improvement, our work provides insights for developing more effective protein multimodal LLMs.
2 Methods
Models: In this paper, we selected four state-of-the-art GDMs specialized in the protein domain, as well as three pre-trained Large Language Models (LLMs). Each model has a distinct embedding dimensionality, adding diversity to our experiments. For GDMs, we used the following models: GAT with an embedding size of 64 [26], ScanNet with an embedding size of 128 [25], GVP with an embedding size of 148 [10], and GearNet with an embedding size of 3072 [31]. All GDMs were pre-trained and loaded directly from their respective official repositories to ensure consistent and stable performance. For LLMs, we selected Gemma2-2B with an embedding size of 2304 [24], LLaMa3.1-8B with an embedding size of 4096 [4], and LLaMa3.1-70B with an embedding size of 8192. All LLMs were loaded directly from Huggingface’s repository.
Dataset: We downloaded protein data from the RCSB PDB database [22]. For this study, we randomly selected 20,000 protein samples. Each protein has two modalities: a text-based description from the FASTA file and a 3D structural graph modality from the PDB file. We randomly split the dataset into 80% for training, 10% for validation, and 10% for testing. A more detailed statistic of the dataset is shown in Appendix A.1.
2.1 Step I: Data Preprocessing
2.1.1 FASTA Preprocessing
Since the selected LLMs are pre-trained on general text data without specific knowledge of proteins, we convert the FASTA file into a clear and detailed text description that the LLMs can understand. This is shown in Figure 1 Data Preprocessing upper part. The process requires stripping unnecessary information and retaining only the essential content. A typical FASTA file contains several key details: protein ID, chains, molecule name, organism, and the amino acid sequence. While we consider most of this information important, the amino acid sequence itself is often lengthy and may introduce noise, so we exclude it from the text description and instead include the sequence length. Additionally, the FASTA file allows us to determine whether a protein consists of a single chain or multiple chains. For single-chain proteins, we use regular expressions to extract relevant details. However, for multi-chain proteins, the complexity increases, and regular expressions alone often fail to capture all information. In such cases, we use GPT-4o to generate accurate descriptions, extracting the chains and corresponding organisms. The GPT-4o prompt is provided in Appendix A.2.
After processing, each protein is described with text similar to: “The protein structure [protein_id] has a sequence length of [number] amino acids. Here is more information: The protein structure [protein_id] involves the following chains: [chains]. The protein is named [protein_name] and is derived from the organism [organism].” We manually validated the GPT-generated descriptions for all multi-chain proteins and found them to be accurate compared to the original FASTA files. Examples of protein descriptions are shown in Appendix A.3.
2.1.2 PDB Preprocessing
[{ ID: ID,
seq: TQDCSFQHSP...’,
coords: [[[74.46, 58.25, -21.65],...],...]}...]
2.2 Step II: Latent Representation Extraction
Once the data has been processed, the next step is to feed it into the respective models for latent representation extraction. This process is shown in Figure 1 “Representation Extraction" part.
2.2.1 LLMs Representation
2.2.2 GDMs Representation
  • GearNet: GearNet encodes protein structure information through a relational graph neural network (GNN) and an edge message-passing mechanism. The input to the model consists of a residue-level relational graph constructed from the 3D structure of the protein, which Gearnet’s code will derived from the input PDB file automatically. The detail is shown in Appendix A.4.1. In the end, to obtain the final protein representation, the features from all hidden layers are concatenated, which results in the final protein representation has the shape $[1,3072]$.
[…]
  • GAT: Previously, we already processed the protein PDB file into the format that GAT supports. Therefore, graph $G$ is fed into a GAT model to compute latent representations. The detail is shown in Appendix A.4.4. In GAT architecture, they use 8 attention heads, with each head computing 8 features, leading to a concatenated output of size $64$ for each node, which results in protein’s final representation size of [1, node_size, 64]. Then we simply perform average pooling to remove the ‘node_size’ dimension. This process yields the latent representation of the protein with size of [1, 64], which will be stored and used later.
2.3 Step III: Representation Alignment
2.3.1 Contrastive Loss Function
To maximize the cosine similarity of positive pairs to 1 and minimize that of negative pairs to 0, we apply a contrastive loss function during training. Specifically, we use a modified version of the InfoNCE loss [20]. Let $g_{i}$ be the projected graph representation for protein $i$, $t_{j}$ be the projected text representation for protein $j$, and $\text{sim}(g_{i},t_{j})$ represent the cosine similarity between $g_{i}$ and $t_{j}$, adjusted to the range $[0,1]$ by using $\frac{\text{sim}(g_{i},t_{j})+1}{2}$. The temperature parameter $\tau$ is set to 0.2 [19], and the batch size $B$ is 32. For each protein $i$, the positive pair similarity is given by $\text{sim}^{+}_{i}=\frac{\text{sim}(g_{i},t_{i})+1}{2}$, and the negative pair similarities are $\text{sim}^{-}_{ij}=\frac{\text{sim}(g_{i},t_{j})+1}{2}$ for all $j\neq i$ within the batch. After calculating these similarities, we apply exponentiation and scale them by the temperature $\tau$:
This overall loss function averages the loss over all proteins in the batch, effectively encouraging high similarity for positive pairs and low similarity for negative pairs.
2.3.2 Implementation Details
3 Results
@@ 3.1 Evaluation Metric
3.2 Model Perspective Analysis
3.2.1 Model Pairs Alignment (RQ1): Which LLM-GDM model pairs demonstrate the best alignment performance?
@@
Figure 2: Alignment Score Details. The figure displays the positive and negative alignment scores for each model pair. The y-axis represents the alignment score, while the x-axis indicates the different LLMs used in the model pairs. Red color correspond to positive scores, and blue color represent negative scores.
@@
Recall that the alignment score of a model pair is determined by its positive and negative scores. In Figure 2, we track these scores, where the x-axis uses ‘untrained’ to indicate the untrained version of each corresponding GDM paired with Gemma2-2B. As shown, the positive and negative scores for untrained models are close to zero, supporting our earlier observation that there is no meaningful alignment without training. However, after training the projection heads, all model pairs show increases in both positive and negative scores, with most model pairs achieving higher positive scores than negative ones, except for those involving GAT, where the negative scores exceed the positive scores. This indicates that even with trained projection heads, alignment remains ineffective for model pairs with GAT, as both positive and negative scores increase at the similar level. In model pairs with GearNet or ScanNet, we observe that the positive scores continue to rise as the LLM size increases, while the negative scores decrease. This explains why alignment scores improve with larger LLMs. However, this raises a question: ‘Why do the negative scores not remain at zero, even though the contrastive loss function is designed to keep them there?’ The answer lies in the way proteins are cataloged and studied. Some proteins have been decomposed and studied multiple times by scientists with different angles, resulting in multiple protein IDs for essentially the same protein. Therefore, even if proteins have different IDs, they may still belong to the same original protein. Additionally, proteins can be homologous, meaning they share structural similarities even if they are technically different proteins. For example, human RNA polymerase II is very similar to yeast RNA polymerase II; despite being distinct proteins with different IDs, they exhibit similarities.
@@
3.2.2 Pearson Correlation (RQ2): Is there a correlation between different model pairs?
We computed the Pearson Correlation Coefficient for all trained model pairs, plus the untrained model pairs that have GearNet on the GDM side. For each protein in the testing set, we first obtain its corresponding GDM and LLM representations. We then calculate the alignment score between these two projected representations for each model pair. This process is repeated across all model pairs, resulting in a list of proteins where each protein is associated with alignment scores from different model pairs. Finally, we analyze these alignment scores to compute the correlation between different model pairs. The results, presented in Figure 3a, reveal several interesting patterns.
Figure 3: a. Pearson Correlation Coefficient. The heatmap shows the Pearson correlation coefficients between different model pairs, with model pairs on the x-axis compared against those on the y-axis. b. GDM Dimension Analysis. It shows the effect of increasing the GearNet embedding dimension (x-axis) on the alignment score (y-axis) for different model pairs.
@@ 3.2.2 Pearson Correlation (RQ2): Is there a correlation between different model pairs?
Figure 4: Protein Perspective Analysis. a. The figure illustrates whether a protein’s amino acid sequence length affects its alignment score across different model pairs. b. The figure illustrates whether rarity of a protein affects its alignment score across different model pairs. c. The figure illustrates whether the number of chains (single or multiple) affects the alignment score across different model pairs, as indicated below each graph.
3.3 Protein Perspective Analysis
3.3.1 RQ3: What types of proteins align well across all model pairs, and which do not?
To analyze the alignment of different proteins, we first define three key properties of proteins: (1) amino acid sequence length, (2) rarity, and (3) single/multiple chains. Within each model pair, we calculate the alignment score between the projected graph and text representations for each protein.
Factor 1: Amino Acid Sequence Length: When analyzing the effect of amino acid sequence length, we conducted experiments on model pairs with Gemma2-2B as LLM side. We categorize the proteins based on their amino acid sequence length. As shown in Figure 4a, the x-axis represents the sequence length, and the y-axis represents the alignment score. Each dot corresponds to a protein with a particular sequence length and its corresponding alignment score from the model pair. We applied simple linear regression to measure the correlation between sequence length and alignment score, which is represented by the trend line. The results indicate that sequence length has no significant impact on alignment performance, as the slope of the regression line is near zero.
@@ 3.3.1 RQ3: What types of proteins align well across all model pairs, and which do not?
We randomly selected 10 popular and 10 rare proteins for analysis and found that rarity significantly affects alignment performance. The details of the random selection and these 20 proteins are provided in Appendix A.5. As shown in Figure 4b, the x-axis represents different model pairs, while the y-axis represents alignment scores. Each dot corresponds to a protein’s alignment score for a specific model pair. Blue dots represent popular proteins, and red dots represent rare proteins. The results clearly show that popular proteins tend to align better, while rare proteins generally show poorer alignment. This observation further supports our proposed challenge in Research Question 1: current protein datasets are often unbalanced, with certain proteins extensively studied by multiple researchers and assigned different IDs despite being the same protein, while others are studied only once. This imbalance explains why popular proteins align better than rare ones, as popular proteins are encountered more frequently during training. Even though most model pairs struggle to align rare proteins, ScanNet- and GearNet-based pairs manage to maintain alignment scores closer to zero, adhering to the behavior expected from the contrastive loss function. In contrast, GVP- and GAT-based pairs exhibit a more scattered distribution of alignment scores, with GAT in particular showing poor alignment for both popular and rare proteins. Another noteworthy finding is that as the size of the LLMs increases, the distribution of proteins’ alignment scores becomes more compact. This is particularly evident in model pairs such as GearNet with Gemma2-2B compared to GearNet with LLaMa3.1-70B, and similarly for ScanNet paired with Gemma2-2B versus ScanNet with LLaMa3.1-70B. In both cases, the alignment scores for popular proteins show a tighter distribution, and this trend also holds for rare proteins. These results are consistent with our previous findings, further emphasizing that model pairs involving ScanNet or GearNet demonstrate better overall alignment and share stronger correlations. Additionally, larger LLMs contribute to better alignment performance, highlighting the importance of embedding size and model capacity in improving multimodal alignment.
@@
3.4 Strategies for Improvement
3.4.1 GDM Dimension Analysis (RQ4): Does increasing the GDM dimension improve alignment performance?
@@
where $y_{i,j}$ is the ground truth label and $\hat{y}_{i,j}$ is the predicted probability for task $j$ and protein $i$. In this way, the model is trained to simultaneously predict multiple binary outcomes, which can capture the presence or absence of different properties or functions in a single sample, such as the functions associated with a particular protein. The dataset used for training is officially provided by TorchDrug [32], a comprehensive machine learning library designed for drug discovery and protein analysis. The dataset consists of three subsets: the training set contains 15,170 samples, the validation set comprises 1,686 samples, and the test set includes 1,860 samples. Each sample in the dataset is annotated with multiple binary labels corresponding to different protein functions. The implementation detail is shown in Appendix A.6.

### v2

# Aligning Large Language Models and Geometric Deep Models for Protein Representation

@@
Aligning Large Language Models and Geometric Deep Models for Protein Representation
Dong Shu Northwestern University, Computer Science Dept, Evanston, IL, 60201, USA Bingbing Duan University of Pittsburgh, Biological Sciences Dept, Pittsburgh, PA, 15260, USA Kai Guo Michigan State University, Computer Science Dept, East Lansing, MI, 48824, USA Kaixiong Zhou North Carolina State University, Electrical and Computer Engineering Dept, Raleigh, NC, 27695, USA Jiliang Tang Michigan State University, Computer Science Dept, East Lansing, MI, 48824, USA Mengnan Du New Jersey Institute of Technology, Data Science Dept, Newark, NJ, 07102, USA
Summary
In this study, we explore the alignment of multimodal representations between LLMs and Geometric Deep Models (GDMs) in the protein domain. We comprehensively evaluate three LLMs with four protein-specialized GDMs. Our work examines alignment factors from both model and protein perspectives, identifying challenges in current alignment methodologies and proposing strategies to improve the alignment process. Experimental results reveal that GDMs incorporating both graph and 3D structural information align better with LLMs, larger LLMs demonstrate improved alignment capabilities, and protein rarity significantly impacts alignment performance. We also find that increasing GDM embedding dimensions, using two-layer projection heads, and fine-tuning LLMs on protein-specific data substantially enhance alignment quality. Lastly, we demonstrate that improved alignment correlates with better downstream performance and reduced hallucination in protein-focused MLLMs.
Introduction
@@
Recently, multimodal large language models (MLLM) such as GPT-4V 1, Gemini 2, Llama-3.2 3, LLaVA-1.5 4, etc., have received increasing attention from the community. A MLLM is an LLM-based model with the ability to receive, reason, and output with multimodal information. This extends beyond just text to potentially include various forms of input and output such as image, audio, video, and other sensory data, allowing the model to process and generate various types of information in an integrated manner. In the realm of protein science, specialized MLLMs such as ProtChatGPT 5, ProteinChat 6 and ProteinGPT 7 have emerged as powerful tools for comprehensive protein analysis. These models can process both protein sequences and structures, integrating this information with vast language understanding capabilities. By combining protein sequence and structure encoders with large language models, these systems can perform complex tasks such as property prediction, structure understanding, and even engage in interactive conversations about protein characteristics. This fusion of modalities allows for a more comprehensive and efficient approach to protein analysis, potentially revolutionizing fields such as drug development and biotechnological advancements.
@@
In these protein-focused MLLMs, latent representation alignment serves as a fundamental technique for integrating different types of protein data. A protein can be represented in multiple modalities - as textual descriptions capturing its function and properties, as graph structures representing amino acid relationships, or as 3D coordinate data describing its spatial configuration. While specialized models exist for each of these representations (like LLMs for text and GDMs for structures), the challenge lies in bridging the gap between these different protein representations to enable comprehensive protein analysis. This is where latent representation alignment plays a crucial role, allowing these diverse protein representations to be mapped into a shared space where they can be meaningfully compared and integrated 5, 6, 7.
@@
Several techniques are commonly used to build multimodal models, each with distinct methods for integrating information from different modalities. In the bio domain, recent models such as Prot2Text 8 integrates Graph Neural Networks (GNNs) 9 and LLMs using an encoder-decoder architecture to generate detailed free-text descriptions of protein functions. MM-StackEns model 10 uses a combination of sequence data processed by a siamese neural network 11 and graph data processed by Graph Attention Networks (GATs) 12 to perform Protein-Protein Interaction (PPI) prediction 13. ProtST framework 14 uses multimodal representation alignment and multimodal mask prediction tasks to align protein sequence representations with biomedical text representations. Among these methods, latent representation alignment remains the most popular due to its simplicity and scalability 15. Protein-focused MLLMs such as ProteinGPT 7 is a multimodal chat system designed for protein analysis, combining protein sequence and structure data with a LLM. This model integrates a protein sequence encoder and structure encoder with a linear projection layer to align diverse embeddings with language model inputs. ProteinChat 6 introduces a ChatGPT-like interface for 3D protein structures, emphasizing interactive protein analysis. Its architecture consists of a GVP 16 block to encode structural information, a projection layer, and a LLM. The protein embeddings from the GVP are aligned with the LLM’s language representations through the projection layer, allowing detailed question-based responses regarding protein 3D structures. ProtChatGPT 5 introduces a multimodal LLM framework that combines protein sequence and structure data using a specialized Protein-Language Pertaining Transformer (PLP-former) to bridge the gap between protein structures and text data. The PLP-former aligns protein data with textual descriptions, while a projection adapter facilitates multimodal integration with the LLM, allowing for accurate question-based protein analysis. By aligning the latent spaces of different modalities, this technique enables seamless cross-modal integration without requiring complex fusion mechanisms or retraining weights for every feature.
@@
Despite the current progress in protein-focused MLLMs, latent representation alignment across different protein modalities remains a significant challenge that requires deeper investigation 7. We aim to address these gaps in understanding and provide insights that could lead to more principled and effective approaches in multimodal protein analysis. In the protein domain, LLMs excel at processing and understanding textual descriptions of proteins, including their functions, properties, and relationships documented in scientific literature 17, 18, 19, 20. Meanwhile, GDMs are specialized neural architectures designed to capture the complex three-dimensional structural information of proteins, including atomic coordinates, bond angles, and spatial relationships between amino acids 21, 22, 23, 24. While both model types offer complementary strengths in protein analysis, effectively aligning their representations remains a significant challenge. Our work explores this alignment challenge through three main areas: I) The model perspective analysis, focusing on alignment performance and correlations between different model pairs. II) The protein perspective analysis, examining which types of proteins align well or poorly across models. III) Strategies for improving alignment, including the impact of GDM representation dimensionality, the effect of projection head complexity, and the benefits of fine-tuning LLMs on protein-specific data. To investigate these areas, we conduct extensive experiments comparing three state-of-the-art LLMs (Gemma2-2B, LLaMa3.1-8B, and LLaMa3.1-70B) with four protein-specialized GDMs (GearNet, GVP, ScanNet, GAT). We analyze the alignment performance of different model pairs, study how protein characteristics affect alignment quality, and evaluate various strategies for enhancing alignment through model architecture modifications and training approaches. An overview of our methodology is provided in Figure 1.
@@
Our exploration of the alignment between LLMs and GDMs in the protein domain has revealed several key findings. We found that GDMs integrating both graph and 3D structural information of proteins tend to align better with LLMs. Additionally, larger LLMs with higher embedding dimensions showed improved alignment performance with the same GDM. Our analysis revealed strong correlations between high-alignment model pairs and other high-alignment pairs. Notably, the rarity of a protein significantly affects the model’s alignment performance, with rare proteins posing greater challenges. Based on our observations, we highlight the challenges and limitations of current protein datasets for representation alignment, particularly due to unequal levels of study across proteins and the presence of homologous relationships among them. In terms of model architecture, we discovered that retraining GDMs to have higher embedding dimensions enhances alignment with LLMs. The complexity of the projection head also plays a crucial role, as increasing the number of layers improves alignment up to a certain threshold, beyond which benefits diminish. Finally, we found that fine-tuning LLMs with protein-specific data can significantly improve alignment with GDMs. By systematically analyzing how different model architectures, protein characteristics, and alignment strategies affect performance, we can provide concrete guidelines for building protein-focused MLLMs that more accurately capture the relationship between protein structure and function. This improved understanding directly translates to better performance in crucial tasks such as protein function prediction, protein-protein interaction analysis, and drug binding site identification.
Results
@@
Model Perspective Analysis
Model Pairs Alignment (RQ1): Which LLM-GDM model pairs demonstrate the best alignment performance?
@@
Recall that the alignment score of a model pair is determined by its positive and negative scores. In Figure 2A, we track these scores, where the x-axis uses ‘untrained’ to indicate the untrained version of each corresponding GDM paired with Gemma2-2B. As shown, the positive and negative scores for untrained models are close to zero, supporting our earlier observation that there is no meaningful alignment without training. However, after training the projection heads, all model pairs show increases in both positive and negative scores, with most model pairs achieving higher positive scores than negative ones, except for those involving GAT, where the negative scores exceed the positive scores. This indicates that even with trained projection heads, alignment remains ineffective for model pairs with GAT, as both positive and negative scores increase at the similar level. In model pairs with GearNet or ScanNet, we observe that the positive scores continue to rise as the LLM size increases, while the negative scores decrease. This explains why alignment scores improve with larger LLMs. However, this raises a question: ‘Why do the negative scores not remain at zero, even though the contrastive loss function is designed to keep them there?’ The answer lies in the way proteins are cataloged and studied. Some proteins have been decomposed and studied multiple times by scientists with different angles, resulting in multiple protein IDs for essentially the same protein. Therefore, even if proteins have different IDs, they may still belong to the same original protein. Additionally, proteins can be homologous, meaning they share structural similarities even if they are technically different proteins. For example, human RNA polymerase II is very similar to yeast RNA polymerase II; despite being distinct proteins with different IDs, they exhibit similarities.
@@
Pearson Correlation (RQ2): Is there a correlation between different model pairs?
We computed the Pearson Correlation Coefficient for all trained model pairs, plus the untrained model pairs that have GearNet on the GDM side. For each protein in the testing set, we first obtain its corresponding GDM and LLM representations. We then calculate the alignment score between these two projected representations for each model pair. This process is repeated across all model pairs, resulting in a list of proteins where each protein is associated with alignment scores from different model pairs. Finally, we analyze these alignment scores to compute the correlation between different model pairs. The results, presented in Figure 2B, reveal several interesting patterns.
Figure 2: Model Perspective Analysis. (A) Alignment Score Details. The figure displays the positive (red dot) and negative (blue dot) alignment scores for each model pair. The y-axis represents the alignment score, while the x-axis indicates the different LLMs used in the model pairs. The difference between the positive and negative score is the model pair alignment performance. (B) Pearson Correlation Coefficient. The heatmap shows the Pearson correlation coefficients between different model pairs, with model pairs on the x-axis compared against those on the y-axis. (C) GDM Dimension Analysis. It shows the effect of increasing the GearNet embedding dimension (x-axis) on the alignment score (y-axis) for different model pairs.
@@
Figure 3: Protein Perspective Analysis. At the top, the figure illustrates whether a protein’s amino acid sequence length affects its alignment score across different model pairs. In the middle, the figure illustrates whether rarity of a protein affects its alignment score across different model pairs. At the bottom, the figure illustrates whether the number of chains (single or multiple) affects the alignment score across different model pairs, as indicated below each graph.
Protein Perspective Analysis
RQ3: What types of proteins align well across all model pairs, and which do not?
To analyze the alignment of different proteins, we define three key properties of proteins: amino acid sequence length, rarity, and single/multiple chains. Within each model pair, we calculate the alignment score between the projected graph and text representations for each protein.
Factor 1: Amino Acid Sequence Length: When analyzing the effect of amino acid sequence length, we conducted experiments on model pairs with Gemma2-2B as LLM side. We categorize the proteins based on their amino acid sequence length. As shown in Figure 3a, the x-axis represents the sequence length, and the y-axis represents the alignment score. Each dot corresponds to a protein with a particular sequence length and its corresponding alignment score from the model pair. We applied simple linear regression to measure the correlation between sequence length and alignment score, which is represented by the trend line. The results indicate that sequence length has no significant impact on alignment, as the slope of the regression line is near zero.
@@
We randomly selected 10 popular and 10 rare proteins for analysis and found that rarity significantly affects alignment performance. The details of the random selection and these 20 proteins are provided in Supplemental Note 5. As shown in Figure 3b, the x-axis represents different model pairs, while the y-axis represents alignment scores. Each dot corresponds to a protein’s alignment score for a specific model pair. Blue dots represent popular proteins, and red dots represent rare proteins. The results clearly show that popular proteins tend to align better, while rare proteins generally show poorer alignment. This observation further supports our proposed challenge in Research Question 1: current protein datasets are often unbalanced, with certain proteins extensively studied by multiple researchers and assigned different IDs despite being the same protein, while others are studied only once. This imbalance explains why popular proteins align better than rare ones, as popular proteins are encountered more frequently during training. Even though most model pairs struggle to align rare proteins, ScanNet- and GearNet-based pairs manage to maintain alignment scores closer to zero, adhering to the behavior expected from the contrastive loss function. In contrast, GVP- and GAT-based pairs exhibit a more scattered distribution of alignment scores, with GAT in particular showing poor alignment for both popular and rare proteins. Another noteworthy finding is that as the size of the LLMs increases, the distribution of proteins’ alignment scores becomes more compact. This is particularly evident in model pairs such as GearNet with Gemma2-2B compared to GearNet with LLaMa3.1-70B, and similarly for ScanNet paired with Gemma2-2B versus ScanNet with LLaMa3.1-70B. In both cases, the alignment scores for popular proteins show a tighter distribution, and this trend also holds for rare proteins. These results are consistent with our previous findings, further emphasizing that model pairs involving ScanNet or GearNet demonstrate better overall alignment and share stronger correlations. Additionally, larger LLMs contribute to better alignment performance, highlighting the importance of embedding size and model capacity in improving multimodal alignment.
@@
Strategies for improvement
GDM Dimension Analysis (RQ4): Does increasing GDM dimension improve alignment?
@@
where $y_{i,j}$ is the ground truth label and $\hat{y}_{i,j}$ is the predicted probability for task $j$ and protein $i$. In this way, the model is trained to simultaneously predict multiple binary outcomes, which can capture the presence or absence of different properties or functions in a single sample, such as the functions associated with a particular protein. The dataset used for training is officially provided by TorchDrug 25, a comprehensive machine learning library designed for drug discovery and protein analysis. The dataset consists of three subsets: the training set contains 15,170 samples, the validation set comprises 1,686 samples, and the test set includes 1,860 samples. Each sample in the dataset is annotated with multiple binary labels corresponding to different protein functions. The implementation detail is shown in Supplemental Note 6.